# Financial sentiment analysis using FinBERT with application in predicting stock movement

### v2

@@ 6 Analysis
From the result we see that our FinBERT sentiment analysis + LSTM model acheives the best training and validation loss. Base BERT + LSTM follows closely with it, which implies that our model has not fully taken advantage of the FinBERT language model since we only use the raw sentiment scores. In this paper, we only look at the sentiment from the titles of news, which contain of relatively small amount information and sentiment. This suggest us also trying to see other informations, for example, social network data like twitter, weibo, etc.
We also see that ARIMA actually capture the long run trend well, whereas our model is able to fit the short run trend which corresponding to the movement in short amount of the time pretty good. For a very short training time compared to other deep learning based models, AIRMA achieves MSE: 0.00975, MAE: 0.07154 and RMSE: 0.09876. As we see from the plot of predicted close price our model tend to under predict the close price in the long run. This suggest that in the future work, we might adopt similar method as in Shi et al. (2022), which preprocesses the data by ARIMA first, and then feed it into the deep neural network.
@@ 7 Conclusion
In this paper, we presented a model that leverages financial sentiment analysis to enhance the prediction of stock market movements. Our approach combined the specialized FinBERT, a transformer-based language model trained on financial corpora, with LSTM networks to interpret and utilize time-series data effectively. We demonstrated that this fusion outperforms traditional models such as ARIMA, as well as LSTM models without sentiment analysis, in capturing the nuances of market fluctuations.
Despite the encouraging results, our study identified several areas for improvement and further research:
@@
    The model exhibits a tendency to underpredict long-term trends. This suggests a potential integration of ARIMA for preprocessing to improve long-range forecasts, as demonstrated by other hybrid approaches in the literature.
@@
    Limitations in the size of the dataset for individual stocks led to concerns about overfitting. Future work could explore the inclusion of a broader array of financial news or social media data to enrich the training set and mitigate this issue.
@@
    While FinBERT’s sentiment analysis showed advantages over base BERT, the margin of improvement was modest. Given the findings in ”Is Domain Adaptation Worth Your Investment? Comparing BERT and FinBERT on Financial Tasks” Peng et al. (2021), it would be prudent to consider further fine-tuning of FinBERT on a larger or augmented dataset, potentially amplifying the benefits of domain-specific modeling.
In conclusion, our sentiment-based stock market prediction model represents a promising step forward in the application of AI to financial forecasting. The interplay between news sentiment and market behavior, while complex, can be harnessed to produce more accurate and reliable predictions. As we continue to refine our model and incorporate a wider spectrum of data sources, we anticipate further improvements in performance and a deeper understanding of the intricate dynamics at play in financial markets.

### v3

@@ 8 Analysis
From the result we see that our FinBERT sentiment analysis + LSTM model acheives the best training and validation loss. Base BERT + LSTM follows closely with it, which implies that our model has not fully taken advantage of the FinBERT language model since we only use the raw sentiment scores from the off-the-shelf ‘ProsusAI/finbert‘ model. Future work should explore fine-tuning FinBERT on task-specific data (e.g., using the NSI proposed in Section 5) or on a larger, more contemporary financial news corpus to potentially enhance its domain-specific capabilities for this prediction task. In this paper, we only look at the sentiment from the titles of news, which contain of relatively small amount information and sentiment. This suggest us also trying to see other informations, for example, social network data like twitter, weibo, etc.
We also see that ARIMA actually capture the long run trend well, whereas our model is able to fit the short run trend which corresponding to the movement in short amount of the time pretty good. For a very short training time compared to other deep learning based models, AIRMA achieves MSE: 0.00975, MAE: 0.07154 and RMSE: 0.09876. As we see from the plot of predicted close price our model tend to under predict the close price in the long run. This suggest that in the future work, we might adopt similar method as in (Shi et al., 2022), which preprocesses the data by ARIMA first, and then feed it into the deep neural network.
@@
8.1 Error Analysis and Limitations
Despite the promising results, our study has several limitations. Firstly, the dataset size for individual stocks, after preprocessing, can be relatively small for training complex deep learning models. For instance, the MSFT dataset reduced to 1440 rows after processing. This limited data availability increases the risk of overfitting, where the model might learn noise specific to the training period rather than generalizable market patterns. This was observed during training where loss dropped significantly within a few epochs.
Secondly, our current sentiment analysis relies solely on news titles. While titles are designed to be informative, they contain a limited amount of textual information compared to the full news article bodies. This might restrict the depth of sentiment that can be extracted.
Thirdly, while FinBERT outperformed base BERT, the improvement margin suggests that the off-the-shelf FinBERT model might not be fully optimized for the specific nuances of our dataset or prediction task.
Future improvements could address these limitations by:
  • Exploring techniques to augment the dataset or incorporate data from multiple related stocks, leveraging inter-stock correlations, to increase the effective training size.
  • Incorporating the full text of news articles or sentiment from other sources like social media (e.g., Twitter) to provide richer contextual information.
  • Investigating further fine-tuning of FinBERT on a larger, more contemporary financial news corpus, or specifically on data labeled with our NSI, to enhance its sentiment extraction capabilities for this task.
@@ 9 Conclusion
In this paper, we presented a model that leverages financial sentiment analysis to enhance the prediction of stock market movements. Our approach combined the specialized FinBERT (‘ProsusAI/finbert‘), a transformer-based language model trained on financial corpora, with LSTM networks to interpret and utilize time-series data effectively. We demonstrated that this fusion outperforms traditional models such as ARIMA, as well as LSTM models without sentiment analysis, in capturing the nuances of market fluctuations.
Despite the encouraging results, our study identified several areas for improvement and further research. Key directions for future work include:
@@
    Hybrid Modeling with ARIMA: The model exhibits a tendency to underpredict long-term trends. Future work could explore integrating ARIMA for preprocessing the time series data to capture long-range dependencies, before feeding the residuals or ARIMA predictions as features into the LSTM module, similar to hybrid approaches suggested by (Shi et al., 2022).
@@
    Data Augmentation and Richer Sentiment Sources: Limitations in the dataset size for individual stocks led to concerns about overfitting. Future work should explore the inclusion of a broader array of financial news, the full body of news articles, or sentiment data from social media platforms to enrich the training set and provide more comprehensive sentiment signals.
@@
    Advanced FinBERT Fine-tuning: While utilizing pre-trained FinBERT showed advantages, the margin of improvement over base BERT was modest. This highlights an opportunity for future work. Building upon findings such as those in ”Is Domain Adaptation Worth Your Investment? Comparing BERT and FinBERT on Financial Tasks” (Peng et al., 2021), further fine-tuning of FinBERT on a larger, more contemporary financial news corpus, or specifically on a dataset labeled with our Numerical Sentiment Index (NSI), could significantly amplify the benefits of domain-specific modeling for this prediction task. Preliminary explorations with alternative labeling strategies for fine-tuning underscored the importance of aligning the fine-tuning task closely with the downstream prediction objective, reinforcing the potential of NSI-based fine-tuning.
  4. 4.
    Exploring Advanced Time-Series Architectures: While LSTM networks are effective, newer transformer-based architectures designed for time-series forecasting, such as TEANet (Zhang et al., 2022), could be investigated as an alternative to or in conjunction with the LSTM module to potentially capture more complex temporal patterns.
In conclusion, our sentiment-based stock market prediction model represents a promising step forward in the application of AI to financial forecasting. The interplay between news sentiment and market behavior, while complex, can be harnessed to produce more accurate and reliable predictions. As we continue to refine our model by exploring these avenues, we anticipate further improvements in performance and a deeper understanding of the intricate dynamics at play in financial markets.